OpenStudy (anonymous):

-5y - 9=-(y - 1) solve please

OpenStudy (uri):

-5y - 9=-(y - 1) Remove the paretheses -5y-9=-y+1 Add 9 both sides -5y-9=-y+1 9 9 ========== @mtdog07 What do you get?

OpenStudy (anonymous):

whats the answer

OpenStudy (uri):

I'm helping you..You need to find the answer :)

OpenStudy (anonymous):

oh, -1 over 2?